OpenStudy (anonymous):

Evaluate (find the value) the following expression: -2-|-18+29| I got 49...the answer to it was -13..anybody care to explain what I did wrong, and how they came up with -13?

OpenStudy (anonymous):

29-18=11 11X-1=-11 -2-11=-13

OpenStudy (anonymous):

thanks..I gotta start writing out each step I guess, instead of doing a bunch of stuff in my head..I tend to miss stuff when I do steps in my head...:P

OpenStudy (anonymous):

those absolute values get weird if you drop a negative

OpenStudy (anonymous):

yeah...and apparently the absolute symbols don't turn the numbers inside into positive numbers, like I was first thinking when I started this section..they're a lot like regular parenthesis in equations.

OpenStudy (anonymous):

oh no they are not they definitely make the number positive but the -1 on the outside reverses that effect do whats on the inside of them first then check for a negative or a square on the outside so you don't get tricked again

OpenStudy (anonymous):

I was thinking the -18 inside the symbols was turned to a positive...

OpenStudy (anonymous):

no you have to do the operation inside the absolute values 1st then change the sign if necessary

OpenStudy (anonymous):

so basically the symbols only turn a number into a positive number, if there's a single negative number inside them, and no numbers on the outside of them?

OpenStudy (anonymous):

it works kind of like this: let a be a real number and b be a real number \[\left| a \pm b \right| \] first you do the operation for instance lets say a+b=c then its the same as \[\left| c \right|=\sqrt{c^2}\]

OpenStudy (anonymous):

? sorry, but you lost me there...

OpenStudy (anonymous):

basically all the rules concerning a+b = whatever...ect...don't really stick with me, from my textbook...it only sticks if I find a pattern when I do the problems, pretty much..

OpenStudy (anonymous):

or the rule is applied to a story problem maybe..